Use the principle of mathematical induction to show that


| sin nx| ≤ n| sin x|


for all n∈ N and for all x ∈ R

1
Expert's answer
2021-11-28T19:07:32-0500

Using the principle of mathematical induction let us show that

"| \sin nx| ≤ n| \sin x|"

for all "n∈ \N" and for all "x ∈ \R."


If "n=1" then the inequality "| \sin x| ≤ | \sin x|" is true for all "x ∈ \R."


Suppose for "n=k" that the inequality "| \sin kx| ≤ k| \sin x|" is true for all "x ∈ \R."


Let us prove the inequality for "n=k+1:"


"|\sin((k+1)x)|=|\sin(kx+x)|=|\sin(kx)\cos x+\cos(kx)\sin x|"


"\le |\sin(kx)|\cdot|\cos x|+|\cos(kx)|\cdot|\sin x|
\le |\sin(kx)|\cdot1+1\cdot|\sin x|"


"\le k|\sin x|+|\sin x|=(k+1)|\sin x|."


We conclude that according to the principle of mathematical induction the inequality "| \sin nx| ≤ n| \sin x|" is true for all "n∈ \N" and for all "x ∈ \R."